1. 500 J of heat is added to 0.620 moles of a monatomic gas. The temperature increases by 15.0 °C. How much work does the gas do as it expands?

a) 116 J

b) 384 J

c) 423 J

d) 616 J

2. 3.00 x 10–3 moles of oxygen gas are sealed in a chamber with a movable piston. The chamber and piston have a radius of 2.50 cm. The mass of the piston is 4.00 kg. What is the pressure of the gas?

a) 8.13 x 104 Pa

b) 1.01 x 104 Pa

c) 1.12 x 104 Pa

d) 1.21 x 104 Pa

3. A gas is compressed in such a way that no heat his transferred. Which of the following is true?

a) the temperature remains constant

b) no work is done

c) the pressure remains constant

d) none of these

4. A gas in a sealed chamber is in good thermal contact with its surroundings. The gas is compressed very very slowly. What type of process is this?

a) isochoric

b) isobaric

c) isothermal

d) adiabatic

5. One mol of diatomic gas is initially at STP. The gas expands isothermally until the volume doubles. How much work did the gas do?

a) 1.57 kJ

b) 1.74 kJ

c) 1.88 kJ

d) not enough info
Answer #1

1) using first law of thermodynamics

Q = dU + W

W = Q - dU

= Q - nCv*dT (Cv for monoatomic gas = 3R/2)

= 500 - 0.62*3*8.314/2 * 15

= 384 J

so option b) 384 J is the correct answer